# Understanding Nonlinear Implicit Bias via Region Counts in Input Space

### v1

@@ A.1 Details on Architectures and Datasets
We conduct experiments on different neural network architectures, including ResNet18 and ResNet34 (he2016deep, ), EfficientNetB0 (tan2019efficientnet, ), SENet18 (hu2018squeeze, ), VGG19 (simonyan2014very, ), MobileNetV2 (sandler2018mobilenetv2, ), ShuffleNetV2 (ma2018shufflenet, ), RegNet200MF (radosavovic2020designing, ), SimpleDLA (yu2018deep, ). We conduct all experiments using NVIDIA RTX 6000 graphics card.
We use CIFAR-10/100 (krizhevsky2009learning, ) and Imagenet-1k (deng2009imagenet, ) as datasets. For CIFAR-10 and CIFAR-100 dataset, each network was trained for $200$ epochs using the Stochastic Gradient Descent (SGD) algorithm with cosine learning rate schedule. We choose 27 combinations of hyperparameters in Table 1, and for each hyperparameter we use 3 random seeds and report the average metrics. For the Imagenet-1k dataset, each network was trained for $50$ epochs with random data crop and random flip. We use the same optimizer and 27 combinations of hyperparameters as in CIFAR-10 and CIFAR-100 experiments. It is worth noting that we make minor adjustments on hyperparameters for certain networks to ensure stable training. For example, in the case of VGG19, the training is unable to converge when the learning rate is set to 0.1; therefore, we adjust it to 0.05.
@@ Appendix D Proof
We first prove two lemmas. See 2
@@
Proof of Theorem 1.
Theorem 1 is a direct consequence of Assumption 2 and the following two lemmas.

### v2

@@ 8 Conclusions and Future Directions
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Details on Architectures and Datasets
We conduct experiments on different neural network architectures, including ResNet18 and ResNet34 (He et al., 2016), EfficientNetB0 (Tan & Le, 2019), SENet18 (Hu et al., 2018), VGG19 (Simonyan & Zisserman, 2014), MobileNetV2 (Sandler et al., 2018), ShuffleNetV2 (Ma et al., 2018), RegNet200MF (Radosavovic et al., 2020), SimpleDLA (Yu et al., 2018). We conduct all experiments using NVIDIA RTX 6000 graphics card.
We use CIFAR-10/100 (Krizhevsky et al., 2009) and Imagenet-1k (Deng et al., 2009) as datasets. For CIFAR-10 and CIFAR-100 dataset, each network was trained for $200$ epochs using the Stochastic Gradient Descent (SGD) algorithm with cosine learning rate schedule. We choose 27 combinations of hyperparameters in Table 1, and for each hyperparameter we use 3 random seeds and report the average metrics. For the Imagenet-1k dataset, each network was trained for $50$ epochs with random data crop and random flip. We use the same optimizer and 27 combinations of hyperparameters as in CIFAR-10 and CIFAR-100 experiments. It is worth noting that we make minor adjustments on hyperparameters for certain networks to ensure stable training. For example, in the case of VGG19, the training is unable to converge when the learning rate is set to 0.1; therefore, we adjust it to 0.05.
@@ Appendix D Proof
We first prove two lemmas. See 6.4
@@
Proof of Theorem 6.3.
Theorem 6.3 is a direct consequence of Assumption 6.2 and the following two lemmas.